# Comparative Analysis of FPGA and GPU Performance for Machine Learning-Based Track Reconstruction at LHCb

### v1

@@ I Introduction
As the complexity of ML models used in high-level triggers for various experiments increases, it becomes crucial to study the fundamental components of these algorithms, such as the basic multilayer perceptron (MLP). Comparing the pros and cons of FPGAs with other processors in various contexts will provide valuable insights into their effectiveness and potential in high-energy physics applications. For this reason, in this work we focus on the first step of the processing chain which involves an MLP. In addition, before the acceleration of the GNN is attempted, the workflow as well as the tools used for it have to be tried and optimised for a simpler model, and the MLP provides the ideal example.
@@
In Section II, we briefly present ETX4VELO, the Graph Neural Network (GNN)-based track-finding pipeline for the VELO, recently implemented in LHCb’s first-level trigger, Allen. It is based on the Exa.TrkX collaboration’s approach which interestingly exhibited a near-linear relationship between the throughput and the input event size [14], contrasting with the quadratic nature of conventional combinatorial algorithms. In the next section, Section III, we present our workflow using HLS4ML for implementing and deploying an MLP model on an FPGA, its throughput, our estimate for the maximum throughput achievable on the Alveo FPGAs as well as its comparison with the GPU implementation.
While the computational performance of the model is strictly evaluated, the model accuracy is not optimized. Future work could include the use of profiling and quantization aware training (QAT) to improve this aspect of the model.

### v2

@@ I Introduction
As the complexity of ML models used in high-level triggers for various experiments increases, it becomes crucial to study the fundamental components of these algorithms, such as the basic multilayer perceptron (MLP). Comparing the pros and cons of FPGAs with other processors in various contexts will provide valuable insights into their effectiveness and potential in high-energy physics applications. For this reason, in this work we focus on the first step of ETX4VELO, the Graph Neural Network (GNN)-based track reconstruction pipeline for the VELO, which involves an MLP. In addition, before the acceleration of the GNN is attempted, the workflow as well as the tools used for it have to be tried and optimized for a simpler model, and the MLP provides the ideal example.
@@
In Section II, we briefly present ETX4VELO, the GNN-based track-finding pipeline for the VELO, recently implemented in LHCb’s first-level trigger, Allen. It is based on the Exa.TrkX collaboration’s approach which interestingly exhibited a near-linear relationship between the throughput and the input event size [14], contrasting with the quadratic nature of conventional combinatorial algorithms. In the next section, Section III, we present our workflow using HLS4ML for implementing and deploying the MLP model involved in the first step of the pipeline on an FPGA, its throughput, our estimate for the maximum throughput achievable on the Alveo FPGAs as well as its comparison with the GPU implementation.